OpenStudy (anonymous):

Solve. x2 + 10x = -21

OpenStudy (anonymous):

IS that x^2 or 2x?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

You need to use the quadratic formula....x^2+10x+21=0

OpenStudy (anonymous):

x^2+10x+21 factor: (x+7)(x+3)=0 x=-7, -3 You don't have to do the quadratic formula in this case because it is factorable.